OpenStudy (anonymous):

evaluate the indefinite integral: [8/xln(4x)]dx

OpenStudy (anonymous):

:D it is 32/(ln( ln (4x) )

OpenStudy (turingtest):

is it\[\int\frac8x\ln(4x)dx\]?

OpenStudy (turingtest):

or\[\int\frac8{x\ln(4x)}dx\]?